QUESTION 5 Sumitomo Banks risk manager has estimated that the DEARs of two of s risk manager has estimated that the DEARs of two of its major assets in its trading portfolio, foreign exchange and bonds, are $150,000 and $250,000, respectively $316,228 ○-$1,106,797 ○-$1,204, 161 -$1,264,911 What is the 10-day VAR of Sumitomos trading portiolio if the correlation among assets is assumed to be -1.0 ○-$100,000
Answer #1

DEAR of a portfolio= [(DEAR1)^2+(DEAR2)^2+(2*DEAR1*DEAR2*correlation)]^0.5

Hence combined DEAR= [(-150,000)^2+(-250,000)^2+(2*-150,000*250,000*-1)]^0.5=$100,000

VaR= DEAR*(days)^0.5=100,000*(10)^0.5=$316,227.77

Hence second option is correct option i.e.$316,228